# Dual Quaternion SE(3) Synchronization with Recovery Guarantees

### v1

@@ 4.3 Efficiency and Accuracy on Real Data
Table 5 visualizes cross-sectional slices of the registered point clouds produced by different methods. Better alignment manifests as sharper, cleaner contours with fewer ghosting artifacts and outliers. Figure 3 shows the evolution of the alignment over DQGPM iterations, where the misregistration decreases monotonically and the point clouds gradually collapse to a consistent configuration. The results for estimation error and CPU time are summarized in Table 4. Overall, DQGPM attains comparable best accuracy with substantially lower runtime than SDR, while maintaining strong performance even when a large fraction of measurements is missing.
@@ A.1 Dual Numbers
We denote the set of dual numbers as $\mathbb{D}$. A dual number $d$ has the form of $d=d_{st}+d_{\mathcal{I}}\epsilon$ with $d_{st}$ and $d_{\mathcal{I}}$ in real number field. We call $d_{st}$ the real part or standard part, and $d_{\mathcal{I}}$ is regarded as the dual part or infinitesimal part. $\epsilon$ is a dual unit satisfying $\epsilon^{2}=0$, which is commutative in multiplication with real numbers, complex numbers, and quaternion numbers. A dual number is called infinitesimal if it only has dual part, otherwise, we call it appreciable.
Addition and multiplication on dual number can be defined as follows: for $d_{i}\in\mathbb{D}$, where $d_{i}=d_{i,st}+d_{i,\mathcal{I}}\epsilon,\ i=1,2$, we have
$d_{1}+d_{2}=(d_{1,st}+d_{2,st})+(d_{1,\mathcal{I}}+d_{2,\mathcal{I}})\epsilon$
$d_{1}d_{2}=d_{1,st}d_{2,st}+(d_{1,st}d_{2,\mathcal{I}}+d_{1,\mathcal{I}}d_{2,st})\epsilon$
Although the above properties show the similarity between dual numbers and complex number field, we should notice that the algebra of dual numbers form a ring but not a field as infinitesimal dual number does not have an inverse element. For the dual number ring, $d=d_{st}+d_{\mathcal{I}}\epsilon\in\mathbb{D}$ with $d_{st}\neq 0$ has an inverse of
$d^{-1}=d_{st}^{-1}(1-d_{\mathcal{I}}d_{st}^{-1}\epsilon),$
@@
$\sqrt{d}=\sqrt{d_{st}}+\frac{d_{\mathcal{I}}}{2\sqrt{d_{st}}}\epsilon$
@@
$|d|=\left\{\begin{aligned} |d_{st}|+\operatorname{sgn}(d_{st})d_{\mathcal{I}}\epsilon,\qquad&\ \rm{if}\ d_{st}\neq 0,\\ |d_{\mathcal{I}}|\epsilon,\qquad&\ \rm{otherwise}.\end{aligned}\right$.
@@ A.3 Dual Quaternion
$q=q_{st}+q_{\mathcal{I}}\epsilon,$
where $q_{st},q_{\mathcal{I}}\in\mathbb{H}$ are referred to as the standard part and the dual part of $q$, respectively. When $q_{st}\neq 0$, the dual quaternion $q$ is called appreciable. Moreover, if both $q_{st}$ and $q_{\mathcal{I}}$ are imaginary quaternions, then $q$ is called an imaginary dual quaternion.
@@
$|q|:=\left\{\begin{aligned} |q_{st}|+{(q_{st}q_{\mathcal{I}}^{*}+q_{\mathcal{I}}q_{st}^{*})\over 2|q_{st}|}\epsilon,&\qquad{\rm if}\ q_{st}\not=0,\\ |q_{\mathcal{I}}|\epsilon,&\qquad{\rm otherwise},\end{aligned}\right$.
@@
$q^{-1}=q_{st}^{-1}-q_{st}^{-1}q_{\mathcal{I}}q_{st}^{-1}\epsilon.$
We denote the set of unit dual quaternions by $\mathrm{UDQ}$. A dual quaternion $q$ is called a unit dual quaternion if $|q|=1$. Any unit dual quaternion is invertible, and its inverse coincides with its conjugate, i.e., $q^{-1}=q^{*}$. Moreover, $q$ is a unit dual quaternion if and only if $q_{st}$ is a unit quaternion and
$q_{st}q_{\mathcal{I}}^{*}+q_{\mathcal{I}}q_{st}^{*}=q_{st}^{*}q_{\mathcal{I}}+q_{\mathcal{I}}^{*}q_{st}=0.$
Consider two dual quaternions $p=p_{st}+p_{\mathcal{I}}\epsilon$ and $q=q_{st}+q_{\mathcal{I}}\epsilon$. Their sum is given by
$p+q=\left(p_{st}+q_{st}\right)+\left(p_{\mathcal{I}}+q_{\mathcal{I}}\right)\epsilon,$
@@
$pq=p_{st}q_{st}+\left(p_{st}q_{\mathcal{I}}+p_{\mathcal{I}}q_{st}\right)\epsilon.$
@@
$\|\bm{x}\|_{2}:=\begin{cases}\sqrt{\sum_{i=1}^{n}|x_{i}|^{2}},&\bm{x}_{st}\neq\bm{0}_{n},\\ \sqrt{\sum_{i=1}^{n}|(x_{i})_{\mathcal{I}}|^{2}}\,\epsilon,&\text{otherwise}.\end{cases}$
@@ Appendix B Representation of SO(3) and SE(3)
$q=q_{0}+\bm{q_{R}}=\cos\frac{\theta}{2}+\bm{\hat{u}}\sin\frac{\theta}{2},$ (B.1)
@@
$x=\left(1+\frac{\epsilon}{2}\bm{t}\right)q,$ (B.2)
@@ Appendix C Proofs in Section 2
Since $y_{11}=\bm{u}_{1}^{*}\bm{C}\bm{u}_{1}=(\bm{x}^{*}\bm{C}\bm{x})/n$, we obtain $\operatorname{tr}(\bm{C}\bm{X}+\bm{X}\bm{C})=2\bm{x}^{*}\bm{C}\bm{x}$. Hence, the argmin of (2.1) equals the argmax of (P), and the optimal values differ only by a factor of $2$.
@@
$\hat{z}$ $=\mathop{\arg\min}\limits_{z\in\mathrm{UDQ}}\|\bm{x}-\hat{\bm{x}}z\|_{2}=\mathop{\arg\max}\limits_{z\in\mathrm{UDQ}}\ \big(z^{*}\hat{\bm{x}}^{*}\bm{x}+(z^{*}\hat{\bm{x}}^{*}\bm{x})^{*}\big).$ (C.1)
@@
and thus (C.1) is equivalent to
$\max_{c,d\in\mathbb{H}}$ $ca+a^{*}c^{*}+\big(cb+b^{*}c^{*}+da+a^{*}d^{*}\big)\epsilon$
$\mathrm{s.t.}$ $dc^{*}+cd^{*}=0,\hskip 18.49988ptc^{*}c=1.$
@@
$c=\frac{a^{*}}{|a|}.$
If $a=0$, then $ca+a^{*}c^{*}=0$ for all unit $c$, hence any $c$ with $c^{*}c=1$ is optimal (the maximizer is not unique). In either case, letting
@@
we have (by the definition of the dual quaternion modulus)
$\max_{z\in\mathrm{UDQ}}\big(z^{*}\hat{\bm{x}}^{*}\bm{x}+(z^{*}\hat{\bm{x}}^{*}\bm{x})^{*}\big)=p+p^{*}=2|\hat{\bm{x}}^{*}\bm{x}|.$
Plugging this into (C.1) yields
$\operatorname{d}^{2}(\hat{\bm{x}},\bm{x})=\min_{z\in\mathrm{UDQ}}\|\bm{x}-\hat{\bm{x}}z\|_{2}^{2}=2n-2|\hat{\bm{x}}^{*}\bm{x}|.$ (C.3)
@@
where the inequality uses the assumption $\bm{x}^{*}\bm{C}\bm{x}\geq\hat{\bm{x}}^{*}\bm{C}\hat{\bm{x}}$. Combining the above display with (C.3)–(C.4) gives
@@
$\operatorname{d}^{2}(\hat{\bm{x}},\bm{x})\leq\frac{2}{n}\cdot 2\sqrt{n}\,\|\bm{\Delta}\|_{{}_{\mathrm{op}}}\operatorname{d}(\hat{\bm{x}},\bm{x})=\frac{4}{\sqrt{n}}\|\bm{\Delta}\|_{{}_{\mathrm{op}}}\operatorname{d}(\hat{\bm{x}},\bm{x}),$
which implies $\operatorname{d}(\hat{\bm{x}},\bm{x})\leq\frac{4}{\sqrt{n}}\|\bm{\Delta}\|_{{}_{\mathrm{op}}}$. This completes the proof.
@@
$y=y_{st}+y_{\mathcal{I}}\epsilon,\qquad z=z_{st}+z_{\mathcal{I}}\epsilon,\qquad u=u_{st}+u_{\mathcal{I}}\epsilon,$
with $y_{st},y_{\mathcal{I}},z_{st},z_{\mathcal{I}},u_{st},u_{\mathcal{I}}\in\mathbb{H}$. Recall that for any dual quaternion $q=q_{st}+q_{\mathcal{I}}\epsilon$,
$|q|^{2}=qq^{*}=|q_{st}|^{2}+2\epsilon\,\mathrm{sc}(q_{st}q_{\mathcal{I}}^{*}),$
@@
Step 1: Expand $|u-z|^{2}$. Since $u,z\in\mathrm{UDQ}$, we have $|u_{st}|=|z_{st}|=1$ and $\mathrm{sc}(u_{st}u_{\mathcal{I}}^{*})=\mathrm{sc}(z_{st}z_{\mathcal{I}}^{*})=0$. Thus,
$|u-z|^{2}$ $=\big|(u_{st}-z_{st})+(u_{\mathcal{I}}-z_{\mathcal{I}})\epsilon\big|^{2}$
$=|u_{st}-z_{st}|^{2}+2\epsilon\,\mathrm{sc}\big((u_{st}-z_{st})(u_{\mathcal{I}}-z_{\mathcal{I}})^{*}\big)$
$=\big(2-2\,\mathrm{sc}(u_{st}z_{st}^{*})\big)-2\epsilon\Big(\mathrm{sc}(u_{st}z_{\mathcal{I}}^{*})+\mathrm{sc}(z_{st}u_{\mathcal{I}}^{*})\Big).$
@@
$4|y-z|^{2}$ $=4\big|(y_{st}-z_{st})+(y_{\mathcal{I}}-z_{\mathcal{I}})\epsilon\big|^{2}$
$=4|y_{st}-z_{st}|^{2}+8\epsilon\,\mathrm{sc}\big((y_{st}-z_{st})(y_{\mathcal{I}}-z_{\mathcal{I}})^{*}\big)$
$=4\big(|y_{st}|^{2}+1-2\,\mathrm{sc}(y_{st}z_{st}^{*})\big)+8\epsilon\Big(\mathrm{sc}(y_{st}y_{\mathcal{I}}^{*})-\mathrm{sc}(y_{st}z_{\mathcal{I}}^{*})-\mathrm{sc}(z_{st}y_{\mathcal{I}}^{*})\Big),$
where we used $\mathrm{sc}(z_{st}z_{\mathcal{I}}^{*})=0$ for $z\in\mathrm{UDQ}$.
@@
Case 1: $y_{st}\neq 0$. By the definition of $\mathcal{N}$, we have $u_{st}=y_{st}/|y_{st}|$. Let $r:=|y_{st}|>0$ and $s:=\mathrm{sc}(y_{st}z_{st}^{*})$. Note that $|z_{st}|=1$ and $|\,\mathrm{sc}(y_{st}z_{st}^{*})\,|\leq|y_{st}z_{st}^{*}|=|y_{st}|=r$, hence $s\in[-r,r]$. Moreover,
$\mathrm{sc}(u_{st}z_{st}^{*})=\mathrm{sc}\Big(\frac{y_{st}}{r}z_{st}^{*}\Big)=\frac{s}{r}.$
@@
$2-2\frac{s}{r}\ \leq\ 4r^{2}+4-8s.$ (C.5)
Indeed, set $t:=s/r\in[-1,1]$. Then (C.5) is equivalent to
@@
If $r\leq\frac{1}{4}$, then $2-8r\geq 0$ and the right-hand side is minimized at t=1t=-1, giving $4r^{2}+2-(2-8r)=4r^{2}+8r\geq 0$. If $r\geq\frac{1}{4}$, then $2-8r\leq 0$ and the right-hand side is minimized at $t=1$, giving $4r^{2}+2+(2-8r)=4(r-1)^{2}\geq 0$. This proves (C.5). Moreover, equality can only occur when $r=1$ and $t=1$, i.e., $|y_{st}|=1$ and $\mathrm{sc}(y_{st}z_{st}^{*})=1$, which implies $y_{st}z_{st}^{*}=1$ and hence $u_{st}=y_{st}=z_{st}$.
Case 2: $y_{st}=0$. In this case $\big(4|y-z|^{2}\big)_{st}=4$. By the definition of $\Pi$, we know $u_{st}=y_{\mathcal{I}}/|y_{\mathcal{I}}|$ and $|u_{st}|=1$. Since $z_{st}$ is also unit, we have $\mathrm{sc}(u_{st}z_{st}^{*})\geq-1$, and thus
$\big(|u-z|^{2}\big)_{st}=2-2\,\mathrm{sc}(u_{st}z_{st}^{*})\leq 4=\big(4|y-z|^{2}\big)_{st}.$
Step 4: Conclude the dual-number inequality. Since $|u-z|^{2}$ and $4|y-z|^{2}$ are dual numbers and the order is lexicographic (standard part first), the standard-part comparison implies $|u-z|^{2}\leq 4|y-z|^{2}$ whenever the standard inequality is strict. In the only equality case from Case 1 (namely $u_{st}=z_{st}$), we have $u-z=(u_{\mathcal{I}}-z_{\mathcal{I}})\epsilon$, hence $|u-z|^{2}=0$, and also $y_{st}=z_{st}$ implies $|y-z|^{2}=0$. Thus the inequality still holds.
@@
Define $\widetilde{\bm{x}}=\Pi(\bm{x})$ as in (2.3). Then

### v2

@@ 4.3 Real Data: Stanford Bunny and 3DMatch
Table 6: Performance on the 3DMatch dataset.
Office1 Office2 Livingroom1 Livingroom2 Redkitchen Hotel3
No. Nodes 53 50 57 47 60 37
Missing 81.67% 83.36% 78.12% 79.86% 70.22% 89.41%
Error_r DQGPM 0.1209 0.0066 7.78e-7 9.00e-6 2.89e-5 3.96e-7
SPEC 1.5195 1.7705 1.3342 2.0904 1.0184 1.2469
SDR 1.5038 1.7616 1.3195 2.0946 1.0082 1.2460
EIG 1.5093 1.7661 1.3238 2.0951 1.0128 1.2469
Error_t DQGPM 0.2454 0.0155 3.12e-6 1.06e-5 4.71e-5 3.87e-7
SPEC 1.8254 1.5662 1.7862 1.6364 0.8466 0.3994
SDR 2.4180 2.0358 2.2108 1.7699 1.1597 0.4336
EIG 1.8860 1.6429 1.7988 1.8343 0.8658 0.4029
Table 3 visualizes cross-sectional slices of the registered point clouds produced by different methods. Better alignment manifests as sharper, cleaner contours with fewer ghosting artifacts and outliers. Figure 3 shows the evolution of the alignment over DQGPM iterations, where the misregistration decreases monotonically and the point clouds gradually collapse to a consistent configuration. The results for estimation error and CPU time are summarized in Table 4. Overall, DQGPM attains comparable accuracy with substantially lower runtime than SDR, while maintaining strong performance even when a large fraction of measurements is missing.
In addition to the Stanford object datasets, we evaluate our method on real indoor scenes from the 3DMatch dataset. Different from the ICP-based construction above, here we directly use the sparse pairwise rigid motions provided in gt.log to build the measurement matrix $C$, and no additional perturbation is added. For each scene, let $G=(V,E)$ denote the graph of point-cloud fragments, where each observed edge $(i,j)\in E$ is associated with a reference relative transformation $\widehat{X}_{ij}\in\mathrm{SE}(3)$. Given the estimated absolute poses $\{X_{i}\}_{i=1}^{n}$, we reconstruct the relative motion on each observed edge as $X_{ij}=X_{i}X_{j}^{-1}$, or equivalently $x_{ij}=x_{i}x_{j}^{*}$ in the dual quaternion representation. Since ground truth is available only on the observed edges, we evaluate directly in the edge space and report
$\mathrm{error}_{r}=\frac{1}{|E|}\sum_{(i,j)\in E}\operatorname{d}_{R}(q_{ij},\widehat{q}_{ij}),$
$\mathrm{error}_{t}=\frac{1}{|E|}\sum_{(i,j)\in E}\operatorname{d}_{T}(t_{ij},\widehat{t}_{ij}),$
where $(q_{ij},t_{ij})$ and $(\widehat{q}_{ij},\widehat{t}_{ij})$ are extracted from $X_{ij}$ and $\widehat{X}_{ij}$, respectively, and $\operatorname{d}_{R},\operatorname{d}_{T}$ are defined as in Section 4.1. The results are summarized in Table 6. The 3D reconstructions in Figure 4 and 5 show that DQGPM produces more accurate and visually consistent alignments than SPEC, SDR, and EIG under highly sparse observations.
(a) DQGPM
(b) SPEC
(c) SDR
(d) EIG
Figure 4: 3D reconstructions of 3DMatch Livingroom2 dataset ($n=47$ and observation missing rate $79.86\%$).
(a) DQGPM
(b) SPEC
(c) SDR
(d) EIG
Figure 5: 3D reconstructions of 3DMatch Office2 dataset ($n=50$ and observation missing rate $83.36\%$).
Acknowledgements
The authors thank Professor Chunfeng Cui and Mr. Hao Yang for their careful reading of an earlier version of the manuscript and for pointing out missing details in the proof of Theorem 3.2. This research is supported in part by the Hong Kong Research Grants Council (RGC) General Research Fund (GRF) project CUHK 14206525.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Dual Numbers
We denote the set of dual numbers as $\mathbb{D}$. A dual number $d$ has the form of $d=d_{\mathrm{st}}+d_{\mathcal{I}}\epsilon$ with $d_{\mathrm{st}}$ and $d_{\mathcal{I}}$ in real number field. We call $d_{\mathrm{st}}$ the real part or standard part, and $d_{\mathcal{I}}$ is regarded as the dual part or infinitesimal part. $\epsilon$ is a dual unit satisfying $\epsilon^{2}=0$, which is commutative in multiplication with real numbers, complex numbers, and quaternion numbers. A dual number is called infinitesimal if it only has dual part, otherwise, we call it appreciable.
Addition and multiplication on dual number can be defined as follows: for $d_{i}\in\mathbb{D}$, where $d_{i}=d_{i,\mathrm{st}}+d_{i,\mathcal{I}}\epsilon,\ i=1,2$, we have
$d_{1}+d_{2}=(d_{1,\mathrm{st}}+d_{2,\mathrm{st}})+(d_{1,\mathcal{I}}+d_{2,\mathcal{I}})\epsilon$
$d_{1}d_{2}=d_{1,\mathrm{st}}d_{2,\mathrm{st}}+(d_{1,\mathrm{st}}d_{2,\mathcal{I}}+d_{1,\mathcal{I}}d_{2,\mathrm{st}})\epsilon$
Although the above properties show the similarity between dual numbers and complex number field, we should notice that the algebra of dual numbers form a ring but not a field as infinitesimal dual number does not have an inverse element. For the dual number ring, $d=d_{\mathrm{st}}+d_{\mathcal{I}}\epsilon\in\mathbb{D}$ with $d_{\mathrm{st}}\neq 0$ has an inverse of
$d^{-1}=d_{\mathrm{st}}^{-1}(1-d_{\mathcal{I}}d_{\mathrm{st}}^{-1}\epsilon),$
@@
$\sqrt{d}=\sqrt{d_{\mathrm{st}}}+\frac{d_{\mathcal{I}}}{2\sqrt{d_{\mathrm{st}}}}\epsilon$
@@
$|d|=\left\{\begin{aligned} |d_{\mathrm{st}}|+\operatorname{sgn}(d_{\mathrm{st}})d_{\mathcal{I}}\epsilon,\quad&\ \rm{if}\ d_{\mathrm{st}}\neq 0,\\ |d_{\mathcal{I}}|\epsilon,\quad&\ \rm{otherwise}.\end{aligned}\right$.
@@ A.3 Dual Quaternion
$q=q_{\mathrm{st}}+q_{\mathcal{I}}\epsilon,$
where $q_{\mathrm{st}},q_{\mathcal{I}}\in\mathbb{H}$ are referred to as the standard part and the dual part of $q$, respectively. When $q_{\mathrm{st}}\neq 0$, the dual quaternion $q$ is called appreciable. Moreover, if both $q_{\mathrm{st}}$ and $q_{\mathcal{I}}$ are imaginary quaternions, then $q$ is called an imaginary dual quaternion.
@@
$|q|:=\left\{\begin{aligned} |q_{\mathrm{st}}|+{(q_{\mathrm{st}}q_{\mathcal{I}}^{*}+q_{\mathcal{I}}q_{\mathrm{st}}^{*})\over 2|q_{\mathrm{st}}|}\epsilon,&\quad{\rm if}\ q_{\mathrm{st}}\not=0,\\ |q_{\mathcal{I}}|\epsilon,&\quad{\rm otherwise},\end{aligned}\right$.
@@
$q^{-1}=q_{\mathrm{st}}^{-1}-q_{\mathrm{st}}^{-1}q_{\mathcal{I}}q_{\mathrm{st}}^{-1}\epsilon.$
We denote the set of unit dual quaternions by $\mathrm{UDQ}$. A dual quaternion $q$ is called a unit dual quaternion if $|q|=1$. Any unit dual quaternion is invertible, and its inverse coincides with its conjugate, i.e., $q^{-1}=q^{*}$. Moreover, $q$ is a unit dual quaternion if and only if $q_{\mathrm{st}}$ is a unit quaternion and
$q_{\mathrm{st}}q_{\mathcal{I}}^{*}+q_{\mathcal{I}}q_{\mathrm{st}}^{*}=q_{\mathrm{st}}^{*}q_{\mathcal{I}}+q_{\mathcal{I}}^{*}q_{\mathrm{st}}=0.$
Consider two dual quaternions $p=p_{\mathrm{st}}+p_{\mathcal{I}}\epsilon$ and $q=q_{\mathrm{st}}+q_{\mathcal{I}}\epsilon$. Their sum is given by
$p+q=\left(p_{\mathrm{st}}+q_{\mathrm{st}}\right)+\left(p_{\mathcal{I}}+q_{\mathcal{I}}\right)\epsilon,$
@@
$pq=p_{\mathrm{st}}q_{\mathrm{st}}+\left(p_{\mathrm{st}}q_{\mathcal{I}}+p_{\mathcal{I}}q_{\mathrm{st}}\right)\epsilon.$
@@
$\|\bm{x}\|_{2}:=\begin{cases}\sqrt{\sum_{i=1}^{n}|x_{i}|^{2}},&\bm{x}_{\mathrm{st}}\neq\bm{0}_{n},\\ \sqrt{\sum_{i=1}^{n}|(x_{i})_{\mathcal{I}}|^{2}}\,\epsilon,&\text{otherwise}.\end{cases}$
@@ Appendix B Representation of SO(3) and SE(3)
$q=q_{0}+\bm{q_{R}}=\cos\frac{\theta}{2}+\bm{\hat{u}}\sin\frac{\theta}{2},$ (11)
@@
$x=\left(1+\frac{\epsilon}{2}\bm{t}\right)q,$ (12)
@@ Appendix C Proofs in Section 2
Since $y_{11}=\bm{u}_{1}^{*}\bm{C}\bm{u}_{1}=(\bm{x}^{*}\bm{C}\bm{x})/n$, we obtain $\operatorname{tr}(\bm{C}\bm{X}+\bm{X}\bm{C})=2\bm{x}^{*}\bm{C}\bm{x}$. Hence, the argmin of (3) equals the argmax of (P), and the optimal values differ only by a factor of $2$.
@@
$\hat{z}$ $=\mathop{\arg\min}\limits_{z\in\mathrm{UDQ}}\|\bm{x}-\hat{\bm{x}}z\|_{2}=\mathop{\arg\max}\limits_{z\in\mathrm{UDQ}}\ \big(z^{*}\hat{\bm{x}}^{*}\bm{x}+(z^{*}\hat{\bm{x}}^{*}\bm{x})^{*}\big).$ (13)
@@
and thus (13) is equivalent to
$\max_{c,d\in\mathbb{H}}$ $ca+a^{*}c^{*}+\big(cb+b^{*}c^{*}+da+a^{*}d^{*}\big)\epsilon$ (15)
$\mathrm{s.t.}$ $dc^{*}+cd^{*}=0,\qquad c^{*}c=1.$
@@
$c=\frac{a^{*}}{|a|},$
and the dual part $d$ for maximizer satisfies $dc^{*}+cd^{*}=0$, which yields a fixed dual part objective for (15).
If $a=0$, the lexicographic maximization reduces to maximizing the dual component in (15). If $b\neq 0$, the dual component is maximized by choosing
$c=\frac{b^{*}}{|b|}.$
For this choice of $c$, any $d$ satisfying the unit-dual-quaternion constraint together with
$db+b^{*}d^{*}=0$
attains the same maximum. If $b=0$, then the dual component is also identically zero, and therefore any unit quaternion $c$ gives an optimal choice of $\hat{z}$. Thus, in all cases, the maximum is attained. Let
@@
we have
$\max_{z\in\mathrm{UDQ}}\bigl(z^{*}\hat{\bm{x}}^{*}\bm{x}+(z^{*}\hat{\bm{x}}^{*}\bm{x})^{*}\bigr)=p+p^{*}=2|\hat{\bm{x}}^{*}\bm{x}|.$
Plugging this into (13) yields
$\operatorname{d}^{2}(\hat{\bm{x}},\bm{x})=\min_{z\in\mathrm{UDQ}}\|\bm{x}-\hat{\bm{x}}z\|_{2}^{2}=2n-2|\hat{\bm{x}}^{*}\bm{x}|.$ (16)
@@
where the inequality uses the assumption $\bm{x}^{*}\bm{C}\bm{x}\geq\hat{\bm{x}}^{*}\bm{C}\hat{\bm{x}}$. Combining the above display with (16)–(17) gives
@@
$\operatorname{d}^{2}(\hat{\bm{x}},\bm{x})\leq\frac{2}{n}\cdot 2\sqrt{n}\,\|\bm{\Delta}\|_{{}_{\mathrm{op}}}\operatorname{d}(\hat{\bm{x}},\bm{x})=\frac{4}{\sqrt{n}}\|\bm{\Delta}\|_{{}_{\mathrm{op}}}\operatorname{d}(\hat{\bm{x}},\bm{x}).$ (18)
Taking standard parts of (18) gives
[…]
Suppose $\|\bm{\Delta}\|_{\mathrm{op},\mathrm{st}}\neq 0$. If $\operatorname{d}_{\mathrm{st}}(\bm{x},\hat{\bm{x}})=0$ and $\|\bm{\Delta}\|_{\mathrm{op},\mathrm{st}}>0$, the desired bound is immediate. Otherwise, $\operatorname{d}_{\rm st}(\bm{x},\hat{\bm{x}})>0$, and multiplying both sides of (18) by $\operatorname{d}(\bm{x},\hat{\bm{x}})^{-1}$ yields
$\operatorname{d}(\bm{x},\hat{\bm{x}})\leq\frac{4}{\sqrt{n}}\|\bm{\Delta}\|_{\mathrm{op}},$
which completes the proof.
@@
$y=y_{\mathrm{st}}+y_{\mathcal{I}}\epsilon,\qquad z=z_{\mathrm{st}}+z_{\mathcal{I}}\epsilon,\qquad u=u_{\mathrm{st}}+u_{\mathcal{I}}\epsilon,$
with $y_{\mathrm{st}},y_{\mathcal{I}},z_{\mathrm{st}},z_{\mathcal{I}},u_{\mathrm{st}},u_{\mathcal{I}}\in\mathbb{H}$. Recall that for any dual quaternion $q=q_{\mathrm{st}}+q_{\mathcal{I}}\epsilon$,
$|q|^{2}=qq^{*}=|q_{\mathrm{st}}|^{2}+2\epsilon\,\mathrm{sc}(q_{\mathrm{st}}q_{\mathcal{I}}^{*}),$
@@
Step 1: Expand $|u-z|^{2}$. Since $u,z\in\mathrm{UDQ}$, we have $|u_{\mathrm{st}}|=|z_{\mathrm{st}}|=1$ and $\mathrm{sc}(u_{\mathrm{st}}u_{\mathcal{I}}^{*})=\mathrm{sc}(z_{\mathrm{st}}z_{\mathcal{I}}^{*})=0$. Thus,
$|u-z|^{2}$ $=\big|(u_{\mathrm{st}}-z_{\mathrm{st}})+(u_{\mathcal{I}}-z_{\mathcal{I}})\epsilon\big|^{2}$
$=|u_{\mathrm{st}}-z_{\mathrm{st}}|^{2}+2\epsilon\,\mathrm{sc}\big((u_{\mathrm{st}}-z_{\mathrm{st}})(u_{\mathcal{I}}-z_{\mathcal{I}})^{*}\big)$
$=\big(2-2\,\mathrm{sc}(u_{\mathrm{st}}z_{\mathrm{st}}^{*})\big)-2\epsilon\Big(\mathrm{sc}(u_{\mathrm{st}}z_{\mathcal{I}}^{*})+\mathrm{sc}(z_{\mathrm{st}}u_{\mathcal{I}}^{*})\Big).$
@@
$4|y-z|^{2}$ $=4\big|(y_{\mathrm{st}}-z_{\mathrm{st}})+(y_{\mathcal{I}}-z_{\mathcal{I}})\epsilon\big|^{2}$
$=4|y_{\mathrm{st}}-z_{\mathrm{st}}|^{2}+8\epsilon\,\mathrm{sc}\big((y_{\mathrm{st}}-z_{\mathrm{st}})(y_{\mathcal{I}}-z_{\mathcal{I}})^{*}\big)$
$=4\big(|y_{\mathrm{st}}|^{2}+1-2\,\mathrm{sc}(y_{\mathrm{st}}z_{\mathrm{st}}^{*})\big)+8\epsilon\Big(\mathrm{sc}(y_{\mathrm{st}}y_{\mathcal{I}}^{*})-\mathrm{sc}(y_{\mathrm{st}}z_{\mathcal{I}}^{*})-\mathrm{sc}(z_{\mathrm{st}}y_{\mathcal{I}}^{*})\Big),$
where we used $\mathrm{sc}(z_{\mathrm{st}}z_{\mathcal{I}}^{*})=0$ for $z\in\mathrm{UDQ}$.
@@
Case 1: $y_{\mathrm{st}}\neq 0$. By the definition of $\mathcal{N}$, we have $u_{\mathrm{st}}=y_{\mathrm{st}}/|y_{\mathrm{st}}|$. Let $r:=|y_{\mathrm{st}}|>0$ and $s:=\mathrm{sc}(y_{\mathrm{st}}z_{\mathrm{st}}^{*})$. Note that $|z_{\mathrm{st}}|=1$ and $|\,\mathrm{sc}(y_{\mathrm{st}}z_{\mathrm{st}}^{*})\,|\leq|y_{\mathrm{st}}z_{\mathrm{st}}^{*}|=|y_{\mathrm{st}}|=r$, hence $s\in[-r,r]$. Moreover,
$\mathrm{sc}(u_{\mathrm{st}}z_{\mathrm{st}}^{*})=\mathrm{sc}\Big(\frac{y_{\mathrm{st}}}{r}z_{\mathrm{st}}^{*}\Big)=\frac{s}{r}.$
@@
$2-2\frac{s}{r}\ \leq\ 4r^{2}+4-8s.$ (19)
Indeed, set $t:=s/r\in[-1,1]$. Then (19) is equivalent to
@@
If $r\leq\frac{1}{4}$, then $2-8r\geq 0$ and the right-hand side is minimized at t=1t=-1, giving $4r^{2}+2-(2-8r)=4r^{2}+8r\geq 0$. If $r\geq\frac{1}{4}$, then $2-8r\leq 0$ and the right-hand side is minimized at $t=1$, giving $4r^{2}+2+(2-8r)=4(r-1)^{2}\geq 0$. This proves (19). Moreover, equality can only occur when $r=1$ and $t=1$, i.e., $|y_{\mathrm{st}}|=1$ and $\mathrm{sc}(y_{\mathrm{st}}z_{\mathrm{st}}^{*})=1$, which implies $y_{\mathrm{st}}z_{\mathrm{st}}^{*}=1$ and hence $u_{\mathrm{st}}=y_{\mathrm{st}}=z_{\mathrm{st}}$.
Case 2: $y_{\mathrm{st}}=0$. In this case $\big(4|y-z|^{2}\big)_{\mathrm{st}}=4$. By the definition of $\Pi$, we know $u_{\mathrm{st}}=y_{\mathcal{I}}/|y_{\mathcal{I}}|$ and $|u_{\mathrm{st}}|=1$. Since $z_{\mathrm{st}}$ is also unit, we have $\mathrm{sc}(u_{\mathrm{st}}z_{\mathrm{st}}^{*})\geq-1$, and thus
$\big(|u-z|^{2}\big)_{\mathrm{st}}=2-2\,\mathrm{sc}(u_{\mathrm{st}}z_{\mathrm{st}}^{*})\leq 4=\big(4|y-z|^{2}\big)_{\mathrm{st}}.$
Step 4: Conclude the dual-number inequality. Since $|u-z|^{2}$ and $4|y-z|^{2}$ are dual numbers and the order is lexicographic (standard part first), the standard-part comparison implies $|u-z|^{2}\leq 4|y-z|^{2}$ whenever the standard inequality is strict. In the only equality case from Case 1 (namely $u_{\mathrm{st}}=z_{\mathrm{st}}$), we have $u-z=(u_{\mathcal{I}}-z_{\mathcal{I}})\epsilon$, hence $|u-z|^{2}=0$, and also $y_{\mathrm{st}}=z_{\mathrm{st}}$ implies $|y-z|^{2}=0$. Thus the inequality still holds.
@@
Define $\widetilde{\bm{x}}=\Pi(\bm{x})$ as in (5). Then